If f(x)+3, g(z) - 8 and h(x) , then Preview Get help: Video License Points possible: 10 Unlimited attempts.
Answer #1

f(x)=x^4+3

g(x)=x-8

h(x)=\sqrt{x}

.

.

f\left(g\left(h\left(x\right)\right)\right)=f\left(g\left(\sqrt{x}\right)\right)

f\left(g\left(h\left(x\right)\right)\right)=f\left(\sqrt{x}-8\right)

{\color{Red} f\left(g\left(h\left(x\right)\right)\right)=\left(\sqrt{x}-8\right)^4+3}

or

further expanding

f\left(g\left(h\left(x\right)\right)\right)=\left(\sqrt{x}-8\right)^2\left(\sqrt{x}-8\right)^2+3

f\left(g\left(h\left(x\right)\right)\right)=(\left(\sqrt{x}\right)^2-2\sqrt{x}\cdot \:8+8^2)+(\left(\sqrt{x}\right)^2-2\sqrt{x}\cdot \:8+8^2)+3

f\left(g\left(h\left(x\right)\right)\right)=(x-16\sqrt{x}+64)(x-16\sqrt{x}+64)+3

f\left(g\left(h\left(x\right)\right)\right)=x^2-16\sqrt{x}*x+64x-16\sqrt{x}*x+256x-1024\sqrt{x}+64x-1024\sqrt{x}+4096+3

{\color{Red} f\left(g\left(h\left(x\right)\right)\right)=x^2-32x\sqrt{x}+384x-2048\sqrt{x}+4096}

.

both answer are right